OpenStudy (anonymous):

Find the surface area of a cone with a radius of 6cm and a height of 14cm. Can you show me the working out?

OpenStudy (johnweldon1993):

I believe this equation for the surface area of a cone is correct \[SA = \pi r^2 + \pi r \sqrt{r^2 + h^2}\] You have 'r' (6) and you have 'h'(14) so solve for the surface area

OpenStudy (johnweldon1993):

Just a reminder....the original equation is \[SA = \pi r ^2 + \pi r s\] where 's' is slant height....but since you do not have the slant height...you can find it via Pythagoras...you are looking for the hypotenuse of that side of the cone...so *after simplifying* you have √(r² + h²)